# AVTENet: Audio-Visual Transformer-based Ensemble Network Exploiting Multiple Experts for Video Deepfake Detection

### v1

@@ Abstract
Forged content shared widely on social media platforms is a major social problem that requires increased regulation and poses new challenges to the research community. The recent proliferation of hyper-realistic deepfake videos has drawn attention to the threat of audio and visual forgeries. Most previous work on detecting AI-generated fake videos only utilizes visual modality or audio modality. While there are some methods in the literature that exploit audio and visual modalities to detect forged videos, they have not been comprehensively evaluated on multi-modal datasets of deepfake videos involving acoustic and visual manipulations. Moreover, these existing methods are mostly based on CNN and suffer from low detection accuracy. Inspired by the recent success of Transformer in various fields, to address the challenges posed by deepfake technology, in this paper, we propose an Audio-Visual Transformer-based Ensemble Network (AVTENet) framework that considers both acoustic manipulation and visual manipulation to achieve effective video forgery detection. Specifically, the proposed model integrates several purely transformer-based variants that capture video, audio, and audio-visual salient cues to reach a consensus in prediction. For evaluation, we use the recently released benchmark multi-modal audio-video FakeAVCeleb dataset. For a detailed analysis, we evaluate AVTENet, its variants, and several existing methods on multiple test sets of the FakeAVCeleb dataset. Experimental results show that our best model outperforms all existing methods and achieves state-of-the-art performance on Testset-I and Testset-II of the FakeAVCeleb dataset.
@@
Deepfake detection, Video forgery detection, Audio-visual feature fusion, Transformer, AVTENet
@@ I Introduction
Figure 1: An example of audio-visual deepfakes. Cyber-attackers aim to create convincing counterfeit videos by manipulating audio and visual streams. Sharing such malicious content can have harmful consequences. Therefore, our proposed method exploits manipulations in both audio and visual streams to effectively detect forged videos.
Advances in social media and machine learning technology have made it easier to generate and spread quickly hyper-realistic counterfeit content, commonly known as Deepfakes. Deepfake refers to forged media content in which the face and/or voice of the source person has been manipulated with the face and/or voice of a target person using deep learning algorithms. The rapid development of modern deep learning technology is driving multimedia processing, achieving incredible results in transforming multimedia content (images, audio, and video). Nowadays, it is a cinch to create or manipulate widely available online media content through various advanced tools and techniques; on top of that, the availability of large-scale datasets grants supplementary potential to deepfake generative models. Models such as Generative Adversarial Networks (GANs) [1] and Variational Autoencoders (VAEs) [2] can create highly realistic synthetic media content that is indistinguishable. Face Swapping, Face Reenactment, Face Editing, and Face Synthesis are different categories of deepfake techniques that help produce convincing and well-crafted facial deepfake media [3], with face swapping being the most common visually manipulated deepfake today. Similarly, Text-to-Speech Synthesis (TTS) and Voice Conversion (VC) are audio deepfake generation techniques that enable deception by cloning an individual’s voice. In addition to deepfake audio generation, automatic speaker verification (SV) systems help recognize the speaker’s identity or identify the speaker in the audio [4, 5]. It has been observed that existing SV systems are easily spoofed by morphing the audio signal. State-of-the-art (SOTA) deepfake generative models have a wide range of beneficial applications [6]; however, deepfakes can have potentially serious adverse effects and can pose a threat when used for malicious purposes, such as revenge porn, cyber-crime, misinformation, political denigration, privacy, legal and ethical issues. Misinformation or disinformation [7] spread through deepfakes can lead to chaos or political instability [8], while cyber-attacks can escalate the threat of access to sensitive information. Prevalent misinformation on online social media leads to the spread of unverified rumors and can lead to social disputes. Therefore, effective deepfake detection methods are urgently needed to mitigate these potential harms, and video deepfake detection has become a focus for media forensics researchers.
@@
At first instance, deepfake content involves uni-modal manipulation in audio or vision. But more recently, multi-modal manipulation has been noticed in generating more realistic deepfake videos. The viral Instagram video of Mark Zuckerberg (Meta’s CEO), as shown in Fig. 1, is the ultimate example of audio-visual deepfakes that combine acoustic and visual forgeries. Detecting such convincing deepfake videos is a crucial but challenging task. To combat the trend of deepfakes, several deep learning-based solutions have been proposed to detect forgeries in videos. However, most existing models only focus on uni-modal detection of video deepfakes, and lack evaluation on multi-modal deepfake datasets. Uni-modal methods limit the performance of detectors to certain scenarios; for example, a video-only detector cannot detect acoustic manipulation, while a detector that relies solely on the audio modality can easily be fooled by visual manipulation. Moreover, the lack of benchmark multi-modal datasets that contain both acoustic and visual forgeries hinders the training of multi-modal video deepfake detectors, as existing video datasets overlook audio deepfakes and multi-modal deepfakes. Therefore, to address the above issues, our study focuses on an audio-visual approach evaluated on a multi-modal deepfake dataset to take a step toward applications in real-world scenarios.
@@
Recently, transformers [9] have emerged as a promising alternative and marked a dominant shift from Convolutional Neural Networks (CNNs) in computer vision tasks due to their ability to model long-range dependencies and capture global context. The sequential processing nature of CNN limits the network’s ability to capture long-range dependencies; however, the attention mechanism in the transformer allows it to model long-range global context and relationships. Existing methods mainly use CNN-based solutions to deal with deepfake detection [10, 11, 12, 13, 14, 15, 16, 17]. These CNN-based methods process input data sequentially, while transformers can process it efficiently by processing the entire input sequence in parallel. Inspired by transformers’ success in various fields, we intend to apply it to the video forgery detection task by leveraging acoustic and visual features. We believe that multiple modalities provide complementary information and stable inferences. Therefore, to fill the above gap, we propose AVTENet, a novel Audio-Visual Transformer-based Ensemble Network that utilizes both acoustic and visual features for effective video forgery detection. The main contributions of our work are summarized as follows:
@@
  • We propose an audio-visual transformer-based ensemble network, AVTENet, demonstrating that audio-only, video-only, and audio-visual transformer-based networks can be effectively combined to improve robustness and detection accuracy. The proposed approach can address scenarios where only a single modality or both modalities in audio vision are manipulated.
@@
  • To understand the strengths and weaknesses of various ensemble approaches, we present four variants of AVTENet and show that the feature fusion approach performs better than majority voting, score fusion, and average score fusion.
@@
  • To evaluate the proposed method, we use the FakeAVCeleb dataset [18], which is a multi-modal audio-visual deepfake dataset. We conduct a comprehensive experimental comparison between the proposed method and existing single-modal and multi-modal methods. The results confirm that our proposed method achieves the highest performance on the test sets of the FakeAVCeleb dataset, outperforming all compared methods.
The remainder of this paper is organized as follows. Sec. II reviews related work. Sec. III illustrates the proposed methodology. Sec. IV outlines the experimental setup and reports the results. Finally, Sec. V provides conclusions and future work.
@@ II-A Video Deepfake Generation
Deepfake technology has become more sophisticated as it uses artificial intelligence (AI) techniques to generate, manipulate, or alter video content, thereby adding significant threats to society, legislation, individuals, community consensus, journalism, and cyber-security. Face2Face [19], Neural Talking-Heads [20], First Order Motion Model for Image Animation [21], and Deep Video Portraits [22] are some of the deepfake generation techniques. Likewise, there exist different types of deepfake videos, including lip-syncing, face-swapping, and voice manipulation. Generative Adversarial Networks (GANs) [23] become a popular deep-learning method to generate realistic images or videos. For instance, in [20] and [24] GAN-based video deepfake generation methods were proposed. The method proposed in [24] generates a dance video of a person from a single image. First, a sequence of human poses is generated and rendered in the video. To generate realistic animated dance videos, this method combines pose estimation, key-point detection, and image synthesis. Similarly, the few-shot learning method proposed in [20] uses a few reference images to generate realistic videos of talking heads. Their method creates a realistic video sequence of a target person speaking by mapping a number of reference images. Another GAN-based method for image-to-image translation was proposed in [25]. To ensure high perceptual qualities of the generated images, the model was trained with a perceptual loss function. The authors evaluated their method on several image generation tasks, such as image colorization, painting, and super-resolution.
@@ II-B Video Deepfake Detection
II-B1 Uni-modal Deepfake Detection
@@
Previous work has mainly focused on a single modality to detect forgeries in videos since synthetic videos generally have less manipulation in the audio track. Some researchers have worked to detect forgeries in videos by exploiting visual artifacts such as inconsistent facial expressions and movements [26, 27], irregular head movements or angles [28, 26], mismatched shadows or lightening [29], distorted or blurred edges around the face or body [30], unnatural eye blinking or movement [31], unnatural background or scenery [32]. The authors of [33] argue that several facial editing techniques exhibit visual artifacts in videos that can be tracked and used to detect video forgeries. Considering this rationale, they demonstrated that even simple visual cues can reveal deepfakes. The authors of [31] used eye blinking, a physiological signal that is not well presented in synthetic videos, to detect fake face videos. Other researchers have exploited emotional inconsistency [26], synchronization between speech and lip movements [11], or motion analysis [34]. A biometrics-based forensics technique was proposed in [35] to detect faceswap deepfakes by exploiting appearance and behavior, especially facial expressions and head movements. MesoNet and MesoInception are two popular models proposed in [12]. Both are compact CNN-based models with a few layers that exploit mesoscopic properties to detect facial tampering in videos. MesoNet analyzes the residual high-frequency content of video frames. XceptionNet [13] is built on depth-wise separable convolutional layers. The authors of [14] and [15] also proposed other CNN-based methods for video forgery detection. The method proposed in [14] employs composite coefficients that effectively scale all dimensions of depth, width, and resolution uniformly. The method proposed in [15] utilizes audio and video streams to extract features and then combines these features to better detect deepfake videos. A transformer network based on an incremental learning strategy is proposed in [36], which utilizes face images and their UV texture generated by a 3D face construction method using a single-face input image to detect synthetic videos. The same authors later proposed a hybrid transformer network to detect forged videos [37]. The transformer network is trained end-to-end using XceptionNet and EfficientNet-B4 as feature extractors.
II-B2 Multi-modal Deepfake Detection
@@
Considering that the auditory context plays a crucial role in visual object perception [38], few recent attempts have addressed the challenge of detecting multi-modal deepfake videos using multiple modalities such as video, audio, and text. For instance, inspired by the Siamese networks and triplet loss, a learning-based approach is proposed in [39], which uses acoustic and visual features of the same video to enhance the information for learning. The authors extracted affective cues and used them to compare the emotions perceived in the two modes to discern whether the video was genuine or manipulated. The authors of [40], [41] and [42] combined audio and visual modalities to develop a video deepfake detector that can handle audio-only, video-only, and audio-visual manipulations. In [40], a promising approach is proposed to exploit the intrinsic synchronization between audio and visual modalities to potentially detect increasingly prevalent deepfakes. In [41], an audio-visual person-of-interest deepfake detector that exploits specific biometrics of people is proposed to cope with the variety of manipulation methods used in synthetic media generation. The high generalizability of the detector is due to its independence from any manipulation techniques and training on videos of real talking faces. Similarly, for audiovisual forgery detection, authors in [42] incorporate audio loss, video loss and audiovisual loss to capture the inconsistency of multi-modality and artifacts from individual modalities. Additionally, ensemble approaches were used in [43] and [10] to improve the accuracy of deep fake detection. In [43], incorporating supplementary textual features has been shown to be beneficial in effectively handling multi-modal fake content detection. The authors built an ensemble network combining uni-modal and cross-modal classifiers to distinguish forged from genuine video clips. On the other hand, the ensemble method in [10] combines CNN-based audio-only, video-only, and audio-visual networks to detect multi-modal deepfakes. Furthermore, the authors of [16] argue that the manipulation of audio or visual flow in video can lead to a lack of harmony between them, such as speech-lip non-synchrony and unnatural facial movements. Therefore, they utilized modality dissonance scores to measure the dissimilarity between two modalities to identify genuine or spoofed videos. Another deep learning method that exploits the audio and visual modalities and considers the importance of speech-lip synchronization for the task of video forgery detection is proposed in [11]. Their audio-visual deepfake detector checks the synchrony between the lip sequence extracted from the video and the synthetic lip sequence generated from the audio track. Similarly, authors in [44] adopted a self-supervised transformer-based approach that utilizes contrastive learning. Their method allows paired acoustic and visual streams to learn mouth motion representations propelling paired representation to be closer and unpaired to stay farther in order to determine audio-visual forgery. Another self-supervised learning approach in [45] exploits an audio-visual temporal synchronization by evaluating consistency between the acoustic stream and faces in a video clip to determine forgery. Although uni-modal methods have been extensively studied in video forgery detection, the underlying correlations implicit between different perception modes have not been fully explored and exploited.
II-C Multi-modal Deepfake Datasets
To the best of our knowledge, DFDC [46] and FakeAVCeleb [18] are the only multi-modal datasets that include both acoustic and visual manipulations in videos.
The DFDC dataset [46] is larger than the FakeAVCeleb dataset, containing 100,000 video clips collected from 3,426 paid subjects. DFDC provides a benchmark for evaluating the performance of deepfake detection models and serves as a crucial resource for advancing research in the field of deepfake detection. It has been released as part of a competition to promote the development of deepfake detection methods. The deepfake videos in this dataset were generated using various deepfake techniques, such as GAN-based methods and non-learning methods. It consists of real and fake videos and considers both acoustic and visual manipulations.
The FakeAVCeleb dataset [18] is a multi-modal audio-visual dataset released after DFDC. The dataset was carefully generated through multiple manipulation methods, taking into account the balance among ethnic backgrounds, gender, and age groups. The real videos of 500 celebrities originating from the Voxceleb2 dataset [47] collected from YouTube form the base set. This base set was further used to generate 20,000 deepfake videos through manipulation techniques, including Faceswap [48], FSGAN [49], Wav2Lip [50], and RTVC (real-time-voice-cloning) [51].
Based on its advantages mentioned above, we use the FakeAVCeleb dataset to evaluate our multi-modal deepfake detector AVTENet. Fig. 2 shows a few samples of each category from the FakeAVCeleb dataset. We do not use the DFDC dataset due to its extreme environmental settings, e.g., some instances are poorly illuminated or overexposed, or sometimes the subject’s face may not be facing the camera. Another reason is that its labels lack separate labels for the acoustic and visual modes in the video.

### v2

# AVTENet: A Human-Cognition-Inspired Audio-Visual Transformer-Based Ensemble Network for Video Deepfake Detection

@@ Abstract
The recent proliferation of hyper-realistic deepfake videos has drawn attention to the threat of audio and visual forgeries. Most previous studies on detecting artificial intelligence-generated fake videos only utilize visual modality or audio modality. While some methods exploit audio and visual modalities to detect forged videos, they have not been comprehensively evaluated on multimodal datasets of deepfake videos involving acoustic and visual manipulations, and are mostly based on convolutional neural networks with low detection accuracy. Considering that human cognition instinctively integrates multisensory information including audio and visual cues to perceive and interpret content and the success of transformer in various fields, this study introduces the audio-visual transformer-based ensemble network (AVTENet). This innovative framework tackles the complexities of deepfake technology by integrating both acoustic and visual manipulations to enhance the accuracy of video forgery detection. Specifically, the proposed model integrates several purely transformer-based variants that capture video, audio, and audio-visual salient cues to reach a consensus in prediction. For evaluation, we use the recently released benchmark multimodal audio-video FakeAVCeleb dataset. For a detailed analysis, we evaluate AVTENet, its variants, and several existing methods on multiple test sets of the FakeAVCeleb dataset. Experimental results show that the proposed model outperforms all existing methods and achieves state-of-the-art performance on Testset-I and Testset-II of the FakeAVCeleb dataset. We also compare AVTENet against humans in detecting video forgery. The results show that AVTENet significantly outperforms humans.
@@
Deepfake detection, audio-visual, AVTENet
@@ I Introduction
Widespread misinformation or disinformation [1] on online social media has fueled the spread of unverified rumors, often escalating into social disputes. Deepfake technology further exacerbates this issue. Deepfake refers to forged media content in which the face and/or voice of the source person are manipulated with the face and/or voice of a target person using deep learning algorithms. Deepfake detection is essential for applications such as digital forensic analysis, content moderation, and protecting public trust in media integrity.
@@
Figure 1: Example of audio-visual deepfakes. Cyber-attackers aim to create convincing counterfeit videos by manipulating audio and visual streams. Sharing such malicious content can have harmful consequences. Therefore, our proposed method exploits manipulations in both audio and visual streams to effectively detect forged videos.
Modern deep learning technology has significantly enhanced the generation or manipulation of media content through various advanced tools and techniques. Generative adversarial networks (GANs) [2] and variational autoencoders (VAEs) [3] can create highly realistic synthetic media content that is indistinguishable from real content. Face swapping, face reenactment, face editing, and face synthesis are various categories of deepfake techniques that help produce convincing and well-crafted facial deepfake media [4]. Initially, deepfake content involved unimodal manipulation of audio or video. But recently, multimodal manipulation has been used to produce more realistic deepfake videos. The viral Instagram video of Mark Zuckerberg (Meta’s CEO), as shown in Fig. 1, is the ultimate example of audio-visual deepfakes that combine acoustic and visual forgeries.
@@
Detecting such convincing deepfake videos is crucial and challenging. To combat the trend of deepfakes, several deep learning-based solutions have been proposed to detect forgeries in videos. Existing deepfake detection methods only focus on unimodal detection of video deepfakes, lacking evaluation on multimodal deepfake datasets. Unimodal methods limit the performance of detectors to certain scenarios; for instance, an audio-only detector (Fig. 2(a)) cannot detect visual manipulation, while a detector that relies solely on the visual modality (Fig. 2(b)) can easily be fooled by acoustic manipulation.
The human brain inherently combines auditory and visual cues to comprehend the world. This multisensory cognitive ability [5, 6] or multimodal data [7] brings greater richness and representation and helps identify inconsistencies or anomalies in deepfake media, as shown in Fig. 2(d). Therefore, by drawing inspiration from human cognitive skills and attention mechanisms, our approach bridges the gap between natural and artificial systems, leveraging auditory and visual sensory information [7] and transformer networks [8] to detect deepfake videos (Fig. 2(c)), and is evaluated on a multimodal dataset to take a step toward applications in real-world scenarios.
Figure 2: Comparison of artificial systems and humans in detecting video forgery.
@@
Recently, transformers [8] have emerged as a promising alternative and marked a dominant shift from convolutional neural networks (CNNs) in computer vision tasks due to their ability to model long-range dependencies and capture global context. The sequential processing nature of CNN limits the ability of the network to capture long-range dependencies; however, the attention mechanism in the transformer allows it to model long-range global context and relationships. Existing methods mainly use CNN-based solutions to deal with deepfake detection [9, 10, 11, 12, 13, 14, 15, 16]. These CNN-based methods process input data sequentially, while transformers can process it efficiently by processing the entire input sequence in parallel. Furthermore, in natural language processing (NLP) and computer vision, many transformer-based foundation models trained with large amounts of data have proven useful for downstream application tasks, such as the video vision transformer (ViViT) [17], the audio spectrogram transformer (AST) [18], and the audio-visual transformer (AV-HuBERT) [19]. Due to the effectiveness of transformers in various fields [20, 21, 22], we intend to apply them to the video forgery detection task by leveraging acoustic and visual features. We believe that multiple modalities provide complementary information and stable inferences [23, 24, 7]. Therefore, to fill the above gap, we propose a novel Audio-Visual Transformer-based Ensemble Network, AVTENet, that utilizes both acoustic and visual features for effective video forgery detection.
Our innovation lies in the application and integration of cutting-edge technologies that were not originally proposed for audio-visual deepfake detection. The main contributions of our work are summarized as follows:
@@
  • We propose an audio-visual transformer-based ensemble network (AVTENet), and demonstrate that its audio-only, video-only, and audio-visual networks respectively based on cutting-edge pre-trained transformer-based foundation models can be effectively combined to improve robustness and detection accuracy. The proposed approach can address scenarios where only a single modality or both modalities in audio vision are manipulated.
@@
  • To evaluate the strengths and weaknesses of various ensemble approaches, we present four variants of AVTENet and show that the feature fusion approach performs better than majority voting, score fusion, and average score fusion.
@@
  • To evaluate the proposed method, we use the FakeAVCeleb dataset [25], which is a multimodal audio-visual deepfake dataset. We conduct a comprehensive experimental comparison between the proposed method and existing unimodal and multimodal methods. The results confirm that our proposed method achieves the highest performance on the test sets of the FakeAVCeleb dataset, outperforming all compared methods.
  • We compare AVTENet’s detection capabilities with human detection capabilities.
The remainder of this paper is organized as follows. Sec. II reviews related work. Sec. III illustrates the proposed methodology. Sec. IV outlines the experimental setup and reports the results. Sec. V discusses the limitations of this study. Finally, Sec VI provides conclusions and future work.
@@ II-A Video Deepfake Generation
Deepfake technology has become more sophisticated as it uses AI techniques to generate, manipulate, or alter video content, thereby adding significant threats to society, legislation, individuals, community consensus, journalism, and cyber-security. Face2Face [26], Neural Talking-Heads [27], First Order Motion Model for Image Animation [28], and Deep Video Portraits [29] are some of the deepfake generation techniques. Likewise, there exist different types of deepfake videos, including lip-syncing, face-swapping, and voice manipulation. GANs [2], [30] have become a popular deep-learning method for generating realistic images or videos. For instance, in [27] and [31] GAN-based video deepfake generation methods were proposed. The method proposed in [31] generates a dance video of a person from a single image. First, a sequence of human poses is generated and rendered in the video. To generate realistic animated dance videos, this method combines pose estimation, key-point detection, and image synthesis. Similarly, the few-shot learning method proposed in [27] uses a few reference images to generate realistic videos of talking heads. Their method creates a realistic video sequence of a target person speaking by mapping several reference images. Another GAN-based method for image-to-image translation was proposed in [32]. To ensure high perceptual qualities of the generated images, the model was trained with a perceptual loss function. The authors evaluated their method on several image generation tasks, such as image colorization, painting, and super-resolution.
@@ II-B Video Deepfake Detection
II-B1 Unimodal Deepfake Detection
@@
Previous work has mainly focused on a single modality to detect forgeries in videos since synthetic videos generally have less manipulation in the audio track (Fig. 2(b)). Some researchers have worked on detecting forgeries in videos by exploiting visual artifacts such as inconsistent facial expressions and movements [33, 34], irregular head movements or angles [35, 33], mismatched shadows or lightening [36], distorted or blurred edges around the face or body [37], unnatural eye blinking or movement [38], unnatural background or scenery [39]. The authors of [40] argued that several facial editing techniques exhibit visual artifacts in videos that can be tracked and used to detect video forgeries. Considering this rationale, they demonstrated that even simple visual cues can reveal deepfakes. The authors of [38] used eye blinking, a physiological signal that is not well presented in synthetic videos, to detect fake face videos. Other researchers have exploited emotional inconsistency [33], synchronization between speech and lip movements [10], or motion analysis [41]. A biometrics-based forensics technique was proposed in [42] to detect faceswap deepfakes by exploiting appearance and behavior, especially facial expressions and head movements. MesoNet and MesoInception are two popular models proposed in [11]. Both are compact CNN-based models with a few layers that exploit mesoscopic properties to detect facial tampering in videos. MesoNet analyzes the residual high-frequency content of video frames. XceptionNet [12] is built on depth-wise separable convolutional layers. The CNN-based method proposed in [13] employs composite coefficients that effectively scale all dimensions of depth, width, and resolution uniformly.
A transformer network based on an incremental learning strategy is proposed in [43], which utilizes face images and their UV texture generated by a 3D face construction method using a single-face input image to detect synthetic videos. Later, the same authors proposed a hybrid transformer network to detect forged videos [44]. The transformer network is trained end-to-end using XceptionNet and EfficientNet-B4 as feature extractors. Obscene video detection [45] can be considered related to the task of fake video detection, where most obscene videos of celebrities are fake.
II-B2 Multimodal Deepfake Detection
@@
Considering that the auditory context is crucial in visual object perception [46], few recent attempts have addressed the challenge of detecting multimodal deepfake videos using multiple modalities (Fig. 2(c)) such as video, audio, and text. The CNN-based method proposed in [14] utilizes audio and video streams to extract features and then combines these features to better detect deepfake videos. A learning-based approach, inspired by Siamese networks and triplet loss, is proposed in [47]. This approach enhances information for learning by using the acoustic and visual features of the same video. The authors extracted affective cues and used them to compare the emotions perceived in the two modes to discern whether the video was genuine or manipulated. The authors of [48], [49], and [50] combined audio and visual modalities to develop a video deepfake detector that can handle audio-only, video-only, and audio-visual manipulations. In [48], a promising approach was proposed to exploit the intrinsic synchronization between audio and visual modalities to potentially detect increasingly prevalent deepfakes. In [49], an audio-visual person-of-interest deepfake detector that exploits specific biometrics of individual was proposed to handle various manipulation methods used in synthetic media generation. The high generalizability of the detector is due to its independence from any manipulation techniques and training on videos of real talking faces. Similarly, for audio-visual forgery detection, the authors of [50] incorporate audio loss, video loss and audio-visual loss to capture the inconsistency of multimodality and artifacts from individual modalities. Additionally, ensemble approaches were used in [9] and [51] to improve the accuracy of deepfake detection. The ensemble method in [9] combines CNN-based audio-only, video-only, and audio-visual networks to detect multimodal deepfakes. In [51], incorporating supplementary textual features was demonstrated to be effective in handling multimodal fake content detection. The authors built an ensemble network combining unimodal and crossmodal classifiers to distinguish forged from genuine video clips. Furthermore, the authors of [15] argued that the manipulation of audio or visual flow in video can lead to a lack of harmony between them, such as speech-lip non-synchrony and unnatural facial movements. Therefore, they utilized modality dissonance scores to measure the dissimilarity between two modalities to identify genuine or spoofed videos. Another deep learning method that exploits the audio and visual modalities and considers the importance of speech-lip synchronization for the task of video forgery detection was proposed in [10]. Their audio-visual deepfake detector checks the synchrony between the lip sequence extracted from the video and the synthetic lip sequence generated from the audio track. Similarly, the authors of [52] adopted a self-supervised transformer-based approach that utilizes contrastive learning. Their method allows paired acoustic and visual streams to learn mouth motion representations propelling paired representation to be closer and unpaired to stay farther to determine audio-visual forgery. Another self-supervised learning approach in [53] exploits an audio-visual temporal synchronization by evaluating consistency between the acoustic stream and faces in a video clip to determine forgery. In addition to the above studies, there are also multimodal methods for other tasks worth referring to. For instance, [54] proposed a two-stage cascade framework based on multimodal data fusion for face anti-spoofing. The authors of [55] proposed a deep CNN– called LieNet to detect deception by combining contact and non-contact signal modalities. The authors of [56] proposed a multimodal feature-fusion-based model called MFFFLD for fingerprint liveness detection. [57] compared the recognition performance and robustness of different multimodal emotion recognition models. Although multimodal approaches have been studied in video forgery detection; more exploration is needed.
@@
II-C Multimodal Deepfake Datasets
To the best of our knowledge, DFDC [58] and FakeAVCeleb [25] are the only multimodal datasets that include both acoustic and visual manipulations in videos.
The DFDC dataset [58] is larger than the FakeAVCeleb dataset, containing 100,000 video clips collected from 3,426 paid subjects. DFDC provides a benchmark for evaluating the performance of deepfake detection models and serves as a crucial resource for advancing research in deepfake detection. It has been released as part of a competition to promote the development of deepfake detection methods. The deepfake videos in this dataset were generated using various deepfake techniques, such as GAN-based methods and nonlearning methods. It comprises real and fake videos and considers both acoustic and visual manipulations.
The FakeAVCeleb dataset [25] is a multimodal audio-visual dataset released after DFDC. The dataset was carefully generated through multiple manipulation methods, considering the balance among ethnic backgrounds, gender, and age groups. The real videos of 500 celebrities originating from the Voxceleb2 dataset [59] collected from YouTube form the base set. This base set was further used to generate 19,500 deepfake videos through manipulation techniques, including Faceswap [60], FSGAN [61], Wav2Lip [62], and RTVC (real-time-voice-cloning) [63].